I would appreciate your help on factoring these polynomials.could you show me how they are suppose to look on my paper?
1)2x^2+2x-8
2)-7a^6+10a^4-14a^2
3)6x^8+12x^6-24x^4+30x^2
I have a hard time knowing how to set them up on paper.Even though we don't have to turn in our homework,I still need to know what I'm doing for the test.My next test is in a couple of days,so I could really benefit from your help.Thank you in advance for giving your time to students like me who are struggling to understand the concepts.
Sincerely,LH

Answer by stanbon(75887)

1)2x^2+2x-8
Notice the common factor of "2".
To find the "other" factor, divide each
term by the common factor. You then get:
= 2(x^2+x-4)
==================
2)-7a^6+10a^4-14a^2
Notice the common factor of "a^2".
Divide each term by a^2 to get:
= a^2(-7a^4+10a^2-14)
OR divide each term by -a^2 to get:
= -a^2(7a^4-10a^2+14)
==================
3)6x^8+12x^6-24x^4+30x^2
Notice the common factor of "6x^2"
Divide each term by 6x^2 to get:
= 6x^2(x^6+2x^4 -4x^2+5)
==============
Cheers,
Stan H.